For this equation, solve for A.
CKA/EN=K^2^/H

Answer by tutorcecilia(2152)

CKA/EN=K^2^/H
(EN/CK)(CKA/EN)=(EN/CK)(K^2^/H)
A=K^2EN/CKH
.
check by plugging the value of A back into the original equation and simplify.